solve by substitution method
y=2x-5
2y-x=2

Answer by Boreal(15235)

y=2x-5
2y-x=2
Substitute (2x-5) for y in the second equation.
2(2x-5) -x =2
distribute the 2
4x-10-x = 2
3x-10 =2
3x=12
x=4
y=2x-5
y=8-5=3
(4,3)
2(3)-4=2 Checks.
Whenever you have y= or x= alone, substitution is usually the easiest way to solve a system of equations.